# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ 3.4 Analyzing the Generalization Ability of the Augmented Graphs via Influence Functions
We used influence functions (Law, 1986; Koh & Liang, 2017; Kong et al., 2021) to understand the impact of augmented data on the model performance on the test set, and thus motivate the use of data augmentation strategy which is specific to the model architecture and the model weights. In Theorem 3.3, we derive a closed-formula for the impact of adding an augmented graph $\mathcal{G}_{n}^{\lambda_{n,m}}$ on the GNN’s performance on a test graph $\mathcal{G}_{k}^{test}$, where the GNN is trained solely on the original training set, without including the augmented graph.
@@
Given a test graph $\mathcal{G}_{k}$ from the test set, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in Eq. 3. The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{n,m}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)$, where $\mathcal{G}_{n}^{\lambda_{n,m}}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{test}$ is given by
@@
$\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}=-\nabla_{\theta}\ell(\mathcal{G}_{k}^{test},\hat{\theta})H_{\hat{\theta}}^{-1}\nabla_{\theta}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta}),$
@@
where $\hat{\theta}_{\epsilon_{n,m}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{n,m}^{\text{aug}}$ denotes the parameters that minimize the upweighted objective function $\mathcal{L}_{n,m}^{\text{aug}}$ and $H_{\hat{\theta}}=\nabla_{\theta}^{2}\mathcal{L}(\hat{\theta})$ is the Hessian Matrix of the loss w.r.t the model parameters.
@@
We provide the proof of Theorem 3.3 in Appendix B. The influence scores are useful for evaluating the effectiveness of the augmented data on each test graph. The strength of influence function theory lies in its ability to analyze the effect of adding augmented data to the training set without actually retraining on this data. As noticed, these influence scores depend not only on the augmented graphs themselves, but also on the model’s weights and architecture. This highlights the need for a graph data augmentation strategy tailored specifically to the GNN backbone in use, as opposed to traditional, techniques like DropNode, DropEdge, and $\mathcal{G}$-Mixup, which are general-purpose methods that can be applied with any GNN architecture.
We can measure the average influence $\mathcal{I}(\mathcal{G}_{n}^{\lambda_{n,m}})$ of a augmented graph $\mathcal{G}_{n}^{\lambda_{n,m}}$ on the whole test set by averaging the derivatives as follows,
$\mathcal{I}(\mathcal{G}_{n}^{\lambda_{n,m}})=\frac{-1}{|\mathcal{D}_{test}|}\sum_{\mathcal{G}_{k}^{test}\in\mathcal{D}_{test}}\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}.$
@@
A negative value of $\mathcal{I}(\mathcal{G}_{n}^{\lambda_{n,m}})$ indicates that adding the augmented data to the training set would increase the prediction loss on the test set, negatively affecting the GNN’s generalization. In contrast, a good augmented graph is one with a postive $\mathcal{I}(\mathcal{G}_{n}^{\lambda_{n,m}})$, indicating improved generalization. In Figure 2, we present the density of the average influence scores of each augmented data on the test set.
@@
4 Empirical Evaluation
4.1 Experimental Setup
Datasets. We evaluate our model on five widely used datasets from the GNN literature, specifically IMDB-BINARY, IMDB-MULTI, PROTEINS, MUTAG, and DD, all sourced from the TUD Benchmark (Morris et al., 2020). These datasets consist of either molecular or social graphs. Detailed statistics for each dataset are provided in Table 8 in Appendix F.
Baselines. We benchmark the performance of our approach against the state-of-the-art graph data augmentation strategies. In particular, we consider the DropNode (You et al., 2020), DropEdge (Rong et al., 2019), SubMix (Yoo et al., 2022), $\mathcal{G}$-Mixup (Han et al., 2022) and GeoMix (Zeng et al., 2024).
Implementation Details. We used the PyTorch Geometric (PyG) open-source library, licensed under MIT (Fey & Lenssen, 2019). The experiments were conducted on an RTX A6000 GPU. For the datasets from the TUD Benchmark, we used a size base split. We utilized two GNN architectures, GIN and GCN, both consisting of two layers with a hidden dimension of 32. The GNN was trained on graph classification tasks for 300 epochs with a learning rate of $10^{-2}$ using the Adam optimizer Kingma & Ba (2014). To model the graph representations of each class, we fit a GMM using the EM algorithm, running for 100 iterations or until the average lower bound gain dropped below $10^{-3}$. The number of Gaussians used in the GMM is provided in Table 9 of Appendix F. In 5 of Appendix D, we also present the performance of using the Variational Bayesian estimation (VB) instead of EM algorithm Tzikas et al. (2008). After generating new graph representations from each GMM, we fine-tuned the post-readout function for 100 epochs, maintaining the same learning rate of $10^{-2}$.
4.2 Experimental Results
On the Generalization of GNN. In Tables 2 and 1, we compare the test accuracy of our data augmentation strategy against baseline methods. Overall, our proposed approach consistently achieves the best or highly competitive performance for most of the datasets. Additionally, we observed that the results of the baseline methods vary depending on the GNN backbone, motivating further investigation using influence functions. As demonstrated in Theorem 3.3, the gradient, and more generally, the model architecture, significantly influence how augmented data impacts the model’s performance on the test set.
@@ 4.2 Experimental Results
Table 1: Classification accuracy ($\pm$ standard deviation) on different benchmark node classification datasets for the data augmentation baselines based on the GCN backbone. The higher the accuracy (in %) the better the model. Highlighted are the first, second best results.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
No Aug. 73.00 (4.94) 47.73 (2.64) 73.92 (5.09) 69.99 (5.35) 69.69 (2.89)
DropEdge 71.70 (5.42) 45.67 (2.46) 73.39 (8.86) 70.07 (3.86) 69.35 (3.37)
DropNode 74.00 (3.44) 43.80 (3.54) 73.89 (8.53) 69.81 (4.61) 69.01 (3.95)
SubMix 72.70 (5.59) 46.00 (2.44) 77.13 (9.69) 67.57 (4.56) 70.11 (4.48)
$\mathcal{G}$-Mixup 72.10 (3.27) 48.33 (3.06) 88.77 (5.71) 65.68 (5.03) 61.20 (3.88)
GeoMix 69.69 (3.37) 49.80 (4.71) 74.39 (7.37) 69.63 (5.37) 68.50 (3.74)
GMM-GDA 71.00 (4.40) 49.82 (4.26) 76.05 (6.47) 70.97 (5.07) 71.90 (2.81)
@@
Table 2: Classification accuracy ($\pm$ standard deviation) on different benchmark node classification datasets for the data augmentation baselines based on the GIN backbone. The higher the accuracy (in %) the better the model. Highlighted are the first, second best results.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
No Aug. 70.30 (3.66) 48.53 (4.05) 83.42 (11.82) 69.54 (3.61) 68.00 (3.18)
DropEdge 70.40 (4.03) 46.80 (3.91) 74.88 (9.62) 68.27 (5.21) 67.82 (4.46)
DropNode 70.30 (3.49) 45.20 (4.24) 75.53 (7.89) 65.40 (4.71) 69.01 (3.95)
SubMix 72.50 (4.98) 48.13 (2.12) 81.90 (9.21) 70.44 (2.58) 68.59 (5.04)
$\mathcal{G}$-Mixup 70.70 (3.10) 47.73 (4.95) 87.77 (7.48) 68.82 (3.48) 63.91 (2.09)
GeoMix 70.60 (4.61) 47.20 (3.75) 81.90 (7.55) 69.80 (5.33) 68.34 (5.30)
GMM-GDA 71.70 (4.24) 49.20 (2.06) 88.83 (5.02) 71.33 (5.04) 68.61 (4.62)
Robustness to Structure Corruption. Besides generalization, we assess the robustness of our data augmentation strategy, following the methodology outlined by (Zeng et al., 2024). Specifically, we test the robustness of data augmentation strategies against graph structure corruption by randomly removing or adding 10% or 20% of the edges in the training set. By corrupting only the training graphs, we introduce a distributional shift between the training and testing datasets. This approach allows us to evaluate GMM-GDA’s ability to generalize well and predict the labels of test graphs, which can be considered OOD examples. The results of these experiments are presented in Table 3 for the IMDB-BINARY, IMDB-MULTI, PROTEINS, and DD datasets. As noted, our data augmentation strategy exhibits the best test accuracy in all cases and improves model robustness against structure corruption.
Influence Functions. In Figure 2, we show the density distribution of the average influence of augmented data sampled using GMM-GDA. For the MUTAG and PROTEINS datasets, we observe that GMM-GDA data augmentation has a positive impact on both GCN and GIN models. In contrast, for the DD dataset, GMM-GDA shows no effect on GIN, while it generates many augmented samples with positive values of the influence scores on GCN, thereby enhancing its performance. These findings are consistent with the empirical results presented in Tables 1 and 2.
Configuration Models. As part of an ablation study, we propose a simple yet effective graph augmentation strategy inspired by Configuration Models (Newman, 2013). As shown in Theorem 3.1, the objective is to control the term $\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathbf{h}_{\mathcal{G}^{\lambda}}-\mathbf{h}_{\mathcal{G}}\right\|\right]$, which can be achieved by regulating the distance between the original and the sampled graph within the input manifold, i.e., $\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right]$. The approach involves generating a sampled version of each training graph while preserving its label by breaking a fraction $q$ of the existing edges (a percentage of the total number of edges $|\mathcal{E}_{n}|$) into half-edges, using a Bernoulli prior distribution $\mathcal{B}(r)$ with probability $r$. This process continues until all half-edges are connected. The strength of this method lies in its simplicity and in preserving the degree distribution, as the degree of each node and the total number of edges in the graph remain unchanged. If the distance norm in the input manifold is the $L_{1}$ distance between adjacency matrix, ||×$|\mathcal{E}|\times r\times r$ is an upper bound of $\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right]$, where $|\mathcal{E}|$ is the average of number of edges in training graphs. The results of this experiment are available in Appendix C. As noticed, the configuration model-based graph augmentation method performs competitively with the baselines and even outperforms them in certain cases. This underscores the importance of Theorem 3.1. When compared to our approach GMM-GDA, the latter gives better results across different datasets and GNN backbones. This difference is primarily due to the configuration model based approach being model-agnostic, whereas GMM-GDA leverages the model’s weights and architecture, as explained in Section 3.4 and supported by Theorem 3.3.
@@
Table 3: Robustness against structure corruption: We present the Classification accuracy ($\pm$ standard deviation). We highlighted the best data augmentation strategy bold. For this experiment, we use the GCN backbone.
@@
Dataset IMDB-BINARY IMDB-MULTI PROTEINS DD IMDB-BINARY IMDB-MULTI PROTEINS DD
DropNode 66.40 (5.51) 44.46 (2.13) 69.18 (4.87) 65.79 (3.23) 64.80 (5.01) 43.06 (2.86) 67.73 (6.43) 64.35 (4.56)
DropEdge 66.70 (5.10) 43.80 (3.11) 69.36 (5.90) 68.42 (4.76) 63.20 (6.30) 41.80 (3.15) 68.10 (5.05) 67.06 (2.53)
SubMix 69.30 (3.76) 46.73 (2.67) 69.80 (4.73) 68.04 (7.64) 63.70 (5.64) 43.73 (3.60) 69.09 (4.58) 59.18 (6.29)
GeoMix 72.20 (5.19) 49.20 (4.31) 70.25 (4.75) 68.00 (3.64) 70.90 (3.85) 48.86 (5.18) 68.36 (6.01) 67.31 (3.91)
$\mathcal{G}$-Mixup 68.30 (5.13) 45.53 (4.12) 61.71 (5.81) 51.26 (8.76) 63.20 (5.54) 44.00 (4.63) 46.63 (5.05) 43.71 (7.12)
NoisyGNN 70.50 (4.71) 40.66 (3.12) 69.45 (4.32) 64.18 (5.71) 63.50 (5.43) 38.66 (4.12) 69.99 (3.78) 63.24 (5.02)
GMM-GDA 72.80 (2.99) 49.36 (4.53) 70.61 (4.30) 68.68 (3.72) 73.10 (3.04) 49.53 (3.54) 70.32 (4.04) 69.01 (3.09)

### v3

@@ 3.4 Analyzing the Generalization Ability of the Augmented Graphs via Influence Functions
We use influence functions (Law, 1986; Koh & Liang, 2017; Kong et al., 2021) to understand the impact of augmented data on the model performance on the test set and thus motivate the use of a data augmentation strategy, which is specific to the model architecture and model weights. In Theorem 3.4, we derive a closed-form formula for the impact of adding an augmented graph $\widetilde{\mathcal{G}}_{n}^{m}$ on the GNN’s performance on a test graph $\mathcal{G}_{k}^{\text{test}}$, where the GNN is trained solely on the training set, without the augmented graph.
@@
Given a test graph $\mathcal{G}_{k}$, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in (1). The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{n,m}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\widetilde{\mathcal{G}}_{n}^{m},\theta)$, where $\widetilde{\mathcal{G}}_{n}^{m}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{\text{test}}$ is given by
@@
$\frac{d\ell(\mathcal{G}_{k}^{\text{test}},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}=-\nabla_{\theta}\ell(\mathcal{G}_{k}^{\text{test}},\hat{\theta})\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}),$
@@
where $\hat{\theta}_{\epsilon_{n,m}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{n,m}^{\text{aug}}$ denotes the parameters that minimize the upweighted objective function $\mathcal{L}_{n,m}^{\text{aug}}$ and $\mathbf{H}_{\hat{\theta}}=\nabla_{\theta}^{2}\mathcal{L}(\hat{\theta})$ is the Hessian matrix of the loss w.r.t. the model parameters.
@@
We provide the proof of Theorem 3.4 in Appendix C. The influence scores are useful for evaluating the effectiveness of the augmented data on each test graph. The strength of influence function theory lies in its ability to analyze the effect of adding augmented data to the training set without actually retraining on this data. As noticed, these influence scores depend not only on the augmented graphs themselves but also on the model’s weights and architecture. This highlights the need for a graph data augmentation strategy tailored specifically to the GNN backbone in use, as opposed to traditional techniques like DropNode, DropEdge, and $\mathcal{G}$-Mixup, which are general-purpose methods that can be applied with any GNN architecture.
Theorem 3.4 is valid for any differentiable loss function. More specifically, if the chosen loss is the cross entropy or the negative log-likelihood, then the Hessian matrix corresponds to the Fisher information matrix (Barshan et al., 2020; Lee et al., 2022). Consequently, the norm of $\mathbf{H}_{\hat{\theta}}^{-1}$, i.e., the inverse of the Hessian matrix, can be bounded above using the Cramér–Rao inequality (Nielsen, 2013). Therefore, a trivial case where the norm of influence scores is zero arises when the gradient of the loss function with respect to the input graphs vanishes. This scenario, for instance, can occur in the DD dataset when using GIN. A detailed analysis of this phenomenon is provided in Section 4. In these cases, data augmentation becomes ineffective, having minimal impact on the GNN’s ability to generalize. We can measure the average influence $\mathcal{I}(\widetilde{\mathcal{G}}_{n}^{m})$ of an augmented graph $\widetilde{\mathcal{G}}_{n}^{m}$ on the test set by averaging the derivatives as follows,
$\mathcal{I}(\widetilde{\mathcal{G}}_{n}^{m})=\frac{-1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathcal{G}_{k}^{\text{test}}\in\mathcal{D}_{\text{test}}}\frac{d\ell(\mathcal{G}_{k}^{\text{test}},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}.$
@@
A negative value of $\mathcal{I}(\widetilde{\mathcal{G}}_{n}^{m})$ indicates that adding the augmented data to the training set would increase the prediction loss on the test set, negatively affecting the GNN’s generalization. In contrast, a good augmented graph is one with a positive $\mathcal{I}(\widetilde{\mathcal{G}}_{n}^{m})$, indicating improved generalization. In Figure 2, we present the density of the average influence scores of each augmented data on the test set.
@@
3.5 Fisher-Guided GMM Augmentation
Using influence scores, we can further improve the generalization of the GNN by filtering candidate augmented representations. The process consists of three key stages. (i) Primary GNN training: The GNN model is first trained on the original training set without incorporating any augmented graphs. (ii) Augmentation and filtering: A pool of candidate augmented graph representations is generated using a data augmentation strategy based on GMMs. Because computing the gradient $\nabla_{\theta}\ell(\mathcal{G}_{k}^{\mathrm{test}},\hat{\theta})$ requires access to ground-truth labels, we evaluate the influence of each candidate augmented graph using the set of validation graph rather than on the unseen test set. This yields a ranking of augmented graphs by their estimated impact on validation performance. During this step, we compute both the gradient and the Hessian only with respect to the post-readout parameters. (iii) Filtering: Finally, we combine a subset of the highest-ranked augmented graphs with the original training set to finetune the post-readout function. This filtering setup aligns perfectly with the assumptions of Theorem 3.4, as we first train the post-readout function without any augmentation, then evaluate each augmentation’s influence, and only afterward retrain the post-readout layer using the selected augmented graphs. Our experiments in Section 4 demonstrate that this training paradigm improves generalization across various datasets and GNN architectures.
4 Experimental Results
@@
Table 1: Classification accuracy ($\pm$ std) on different benchmark graph classification datasets for the data augmentation baselines based on the GCN backbone. The higher the accuracy (in %) the better the model. Highlighted are the first, second best results.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
No Aug. $73.00{\scriptstyle\pm 4.94}$ $47.73{\scriptstyle\pm 2.64}$ $73.92{\scriptstyle\pm 5.09}$ $69.99{\scriptstyle\pm 5.35}$ $69.69{\scriptstyle\pm 2.89}$
DropEdge $71.70{\scriptstyle\pm 5.42}$ $45.67{\scriptstyle\pm 2.46}$ $73.39{\scriptstyle\pm 8.86}$ $70.07{\scriptstyle\pm 3.86}$ $69.35{\scriptstyle\pm 3.37}$
DropNode $\mathbf{74.00{\scriptstyle\pm 3.44}}$ $43.80{\scriptstyle\pm 3.54}$ $73.89{\scriptstyle\pm 8.53}$ $69.81{\scriptstyle\pm 4.61}$ $69.01{\scriptstyle\pm 3.95}$
SubMix $\underline{72.70{\scriptstyle\pm 5.59}}$ $46.00{\scriptstyle\pm 2.44}$ $\underline{77.13{\scriptstyle\pm 9.69}}$ $67.57{\scriptstyle\pm 4.56}$ $70.11{\scriptstyle\pm 4.48}$
$\mathcal{G}$-Mixup $72.10{\scriptstyle\pm 3.27}$ $48.33{\scriptstyle\pm 3.06}$ $\mathbf{88.77{\scriptstyle\pm 5.71}}$ $65.68{\scriptstyle\pm 5.03}$ $61.20{\scriptstyle\pm 3.88}$
GeoMix $69.69{\scriptstyle\pm 3.37}$ $\underline{49.80{\scriptstyle\pm 4.71}}$ $74.39{\scriptstyle\pm 7.37}$ $69.63{\scriptstyle\pm 5.37}$ $68.50{\scriptstyle\pm 3.74}$
GRATIN $71.00{\scriptstyle\pm 4.40}$ $\mathbf{49.82{\scriptstyle\pm 4.26}}$ $76.05{\scriptstyle\pm 6.74}$ $\mathbf{70.97{\scriptstyle\pm 5.07}}$ $\mathbf{71.90{\scriptstyle\pm 2.81}}$
@@
Table 2: Classification accuracy ($\pm$ std) on different benchmark graph classification datasets for the data augmentation baselines based on the GIN backbone. The higher the accuracy (in %) the better the model. Highlighted are the first, second best results.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
No Aug. $70.30{\scriptstyle\pm 3.66}$ $\underline{48.53{\scriptstyle\pm 4.05}}$ $83.42{\scriptstyle\pm 2.12}$ $69.54{\scriptstyle\pm 3.61}$ $68.00{\scriptstyle\pm 3.18}$
DropEdge $70.40{\scriptstyle\pm 4.03}$ $46.80{\scriptstyle\pm 3.91}$ $74.88{\scriptstyle\pm 9.62}$ $68.27{\scriptstyle\pm 5.21}$ $67.82{\scriptstyle\pm 4.46}$
DropNode $70.30{\scriptstyle\pm 3.49}$ $45.20{\scriptstyle\pm 4.24}$ $75.53{\scriptstyle\pm 7.89}$ $65.40{\scriptstyle\pm 4.71}$ $\mathbf{69.01{\scriptstyle\pm 3.95}}$
SubMix $\mathbf{72.50{\scriptstyle\pm 4.98}}$ $48.13{\scriptstyle\pm 2.12}$ $81.90{\scriptstyle\pm 9.21}$ $\underline{70.44{\scriptstyle\pm 2.58}}$ $68.59{\scriptstyle\pm 5.04}$
$\mathcal{G}$-Mixup $70.70{\scriptstyle\pm 3.10}$ $47.73{\scriptstyle\pm 4.95}$ $87.77{\scriptstyle\pm 7.48}$ $68.82{\scriptstyle\pm 3.48}$ $63.91{\scriptstyle\pm 2.09}$
GeoMix $70.60{\scriptstyle\pm 4.61}$ $47.20{\scriptstyle\pm 3.75}$ $81.90{\scriptstyle\pm 7.55}$ $69.80{\scriptstyle\pm 5.33}$ $68.34{\scriptstyle\pm 5.30}$
GRATIN $71.70{\scriptstyle\pm 4.24}$ $\mathbf{49.20{\scriptstyle\pm 2.06}}$ $\mathbf{88.83{\scriptstyle\pm 5.02}}$ $\mathbf{71.33{\scriptstyle\pm 5.04}}$ $\underline{68.61{\scriptstyle\pm 4.62}}$
@@
Table 3: Robustness against structure corruption: We present the Classification accuracy ($\pm$ std). The higher the accuracy (in %) the better the model. We highlighted the best data augmentation strategy bold. For this experiment, we use the GCN backbone.
@@
Dataset IMDB-BIN IMDB-MUL PROTEINS DD IMDB-BIN IMDB-MUL PROTEINS DD
DropNode $66.40{\scriptstyle\pm 5.51}$ $44.46{\scriptstyle\pm 2.13}$ $69.18{\scriptstyle\pm 4.87}$ $65.79{\scriptstyle\pm 3.23}$ $64.80{\scriptstyle\pm 5.01}$ $43.06{\scriptstyle\pm 2.86}$ $67.73{\scriptstyle\pm 6.43}$ $64.35{\scriptstyle\pm 4.56}$
DropEdge $66.70{\scriptstyle\pm 5.10}$ $43.80{\scriptstyle\pm 3.11}$ $69.36{\scriptstyle\pm 5.90}$ $68.42{\scriptstyle\pm 4.76}$ $63.20{\scriptstyle\pm 6.30}$ $41.80{\scriptstyle\pm 3.15}$ $68.10{\scriptstyle\pm 5.05}$ $67.06{\scriptstyle\pm 2.53}$
SubMix $69.30{\scriptstyle\pm 3.76}$ $46.73{\scriptstyle\pm 2.67}$ $69.80{\scriptstyle\pm 4.73}$ $68.04{\scriptstyle\pm 7.64}$ $63.70{\scriptstyle\pm 5.64}$ $43.73{\scriptstyle\pm 3.60}$ $69.09{\scriptstyle\pm 4.58}$ $59.18{\scriptstyle\pm 6.29}$
GeoMix $72.20{\scriptstyle\pm 5.19}$ $49.20{\scriptstyle\pm 4.31}$ $70.25{\scriptstyle\pm 4.75}$ $68.00{\scriptstyle\pm 3.64}$ $70.90{\scriptstyle\pm 3.85}$ $48.86{\scriptstyle\pm 5.18}$ $68.36{\scriptstyle\pm 6.01}$ $67.31{\scriptstyle\pm 3.91}$
$\mathcal{G}$-Mixup $68.30{\scriptstyle\pm 5.13}$ $45.53{\scriptstyle\pm 4.12}$ $61.71{\scriptstyle\pm 5.81}$ $51.26{\scriptstyle\pm 8.76}$ $63.20{\scriptstyle\pm 5.54}$ $44.00{\scriptstyle\pm 4.63}$ $46.63{\scriptstyle\pm 5.05}$ $43.71{\scriptstyle\pm 7.12}$
NoisyGNN $70.50{\scriptstyle\pm 4.71}$ $40.66{\scriptstyle\pm 3.12}$ $69.45{\scriptstyle\pm 4.32}$ $64.18{\scriptstyle\pm 5.71}$ $63.50{\scriptstyle\pm 5.43}$ $38.66{\scriptstyle\pm 4.12}$ $69.99{\scriptstyle\pm 3.78}$ $63.24{\scriptstyle\pm 5.02}$
GRATIN $\mathbf{72.80{\scriptstyle\pm 2.99}}$ $\mathbf{49.36{\scriptstyle\pm 4.53}}$ $\mathbf{70.61{\scriptstyle\pm 4.30}}$ $\mathbf{68.68{\scriptstyle\pm 3.72}}$ $\mathbf{73.10{\scriptstyle\pm 3.04}}$ $\mathbf{49.53{\scriptstyle\pm 3.54}}$ $\mathbf{70.32{\scriptstyle\pm 4.04}}$ $\mathbf{69.01{\scriptstyle\pm 3.09}}$
In this section, we present our results and analysis. Our experimental setup is described in Appendix K.
On the Generalization of GNNs. In Tables 1 and 2, we compare the test accuracy of our data augmentation strategy against baseline methods. Additional results for the same experiment on larger datasets can be found in Appendix L. We trained all baseline models using the same train/validation/test splits, GNN architectures, and hyperparameters to ensure a fair comparison. It is worth noting that the baselines exhibit high standard deviations, which is a common characteristic in graph classification tasks. Unlike node classification, graph classification is known to have a larger variance in performance metrics (Errica et al., 2020; Duval & Malliaros, 2022). Overall, our proposed approach consistently achieves the best or highly competitive performance for most of the datasets.
Additionally, we observed that the results of the baseline methods vary depending on the GNN backbone, motivating further investigation using influence functions. As demonstrated in Theorem 3.4, the gradient, and more generally, the model architecture, significantly influence how augmented data impacts the model’s performance on the test set.
Robustness to Structure Corruption. Besides generalization, we assess the robustness of our data augmentation strategy, following the methodology outlined by (Zeng et al., 2024). Specifically, we test the robustness of data augmentation strategies against graph structure corruption by randomly removing or adding 10% or 20% of the edges in the training set. By corrupting only the training graphs, we introduce a distributional shift between the training and testing datasets. This approach allows us to evaluate GRATIN’s ability to generalize well and predict the labels of test graphs, which can be considered OOD examples. The results of these experiments are presented in Table 3 for the IMDB-BIN, IMDB-MUL, PROTEINS, and DD datasets. As noted, our data augmentation strategy exhibits the best test accuracy in all cases and improves model robustness against structure corruption.
Influence Functions. In Figure 2, we show the density distribution of the average influence of augmented data sampled using GRATIN. These findings are consistent with the empirical results presented in Tables 1 and 2. For the MUTAG and PROTEINS datasets, we observe that GRATIN’s data augmentation has a positive impact on both GCN and GIN models. In contrast, for the DD dataset, GRATIN shows no effect on GIN, while it generates many augmented samples with positive values of the influence scores on GCN, thereby enhancing its performance. This behavior is consistent with the baselines, as most graph data augmentation strategies tend to enhance test accuracy more significantly for GCN than for GIN when applied to DD. This is an interesting phenomenon and worthy of deeper analysis. For the DD dataset, when using the GIN model at inference, we observe Softmax saturation, where the predicted class probabilities approach extreme values (close to 00 or $1$), c.f. Appendix M. We suspect this saturation to happen due to the large average number of nodes in DD and the fact that GIN does not normalize the node representations, which may lead to a graph representation with large norms. This saturation results in the model making predictions with very high confidence. Consequently, the gradient of the loss function with respect to the input graphs eventually converges to 00. In such cases, the influence scores become negligible, as explained in Section 3.4.
Fisher Based Filtering. Figure 3 illustrates the impact of removing augmented representations on test accuracy in the Fisher-Guided GMM Augmentation experiment. The results presented are from a single training run. At the beginning, removing augmented graphs with low or negative influence scores improves generalization. The highest test accuracy is reached when a significant portion of low-quality augmentations has been removed while retaining high-influence ones. This indicates that a well-selected augmentation subset enhances model performance. As more augmentations are removed, the overall diversity of the training set decreases. Since data augmentation generally helps the model generalize better, excessive removal reduces its effectiveness. At 100% removal, augmentation is entirely disabled, meaning the model is trained only on the original dataset, i.e., the reference case, leading to a significant drop in accuracy.
Figure 3: Effect of filtering augmented representations on test accuracy.
Configuration Models. As part of an ablation study, we propose a simple yet effective graph augmentation strategy inspired by configuration models (Newman, 2013). As shown in Theorem 3.1, the objective is to control the term $\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\|\mathbf{h}_{\widetilde{\mathcal{G}}}-\mathbf{h}_{\mathcal{G}}\|\right]$, which can be achieved by regulating the distance between the original and the sampled graph within the input manifold, i.e., $\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\|\widetilde{\mathcal{G}}-\mathcal{G}\|\right]$. The approach involves generating a sampled version of each training graph by randomly breaking existing edges into half-edges with probability $r$ and then randomly connecting half-edges until all edges are connected. The strength of this method lies in its simplicity and in preserving the degree distribution. If the distance norm is the $L_{1}$ distance between adjacency matrices, $|\mathcal{E}|r^{2}$ is an upper bound of $\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\|\widetilde{\mathcal{G}}-\mathcal{G}\|\right]$, where $|\mathcal{E}|$ is the average number of edges. The results of this experiment are available in Appendix E.